### v1

# thanks: deceasedthanks: deceased

Abstract
We conducted measurements of $K^{-}+{{}^{3}{\rm He}}\to\pi Y\!N+N^{\prime}$ reactions using $1~{}{\rm GeV}/c$ $K^{-}$-beam, aiming to understand the broad decay width of $\bar{K}\!N\!N$ ($\approx$ twice as broad as that of $\Lambda(1405)$ considered to be the $\bar{K}\!N$ quasi-bound state). We successfully reproduced distributions of the $\pi Y\!N$ invariant mass and momentum transfer of $\pi Y\!N$ by employing model fitting functions for $\bar{K}\!N\!N$ formation and quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) processes. The model can describe the experimental data quite well, and four $\bar{K}NN\to\pi Y\!N$ cross-sections were obtained. The result suggests that the mesonic decay is the dominant decay branch of $\bar{K}\!N\!N$. It was also suggested that $\Gamma_{\pi\Lambda N}\sim\Gamma_{\pi\Sigma N}$, which indicates a significant contribution of the $I_{\bar{K}\!N}=1$ absorption channel to the $\bar{K}\!N\!N$ decay as well as the $I_{\bar{K}\!N}=0$ makes the $\bar{K}\!N\!N$ state $\approx$ twice as unstable as $\Lambda$(1405).
preprint: APS/123-QED
I Introduction
Quasi-bound systems of antikaon ($\bar{K}$) and nucleus, so-called kaonic nuclei, represents an exotic nuclear state wherein a meson exists as a constituent particle inside nuclei. The existence of a kaonic nucleus is predicted as a natural extension of the interpretation of the $\Lambda(1405)$, which is not considered a simple three-quark state but rather a $\bar{K}$-$N$ hadronic molecule attributed to the strongly attractive $\bar{K}N$ interaction in the isospin $I_{\bar{K}N}=0$ channel. Properties of kaonic nuclei, such as binding energies, widths, and branching ratios, nor even confirmation of the existence of kaonic nuclei, are essential to construct a precise model of the $\bar{K}N$ interaction. Understanding these properties allows us to determine the energy dependence of the $\bar{K}\!N$ interaction or the $\bar{K}$-absorption strength to nucleon(s) in the energy region below the $\bar{K}N$ mass threshold. This energy region cannot be directly probed by $\bar{K}N$ scattering or $K^{-}$-atom experiments.
Among kaonic nuclei, the lightest system, $\bar{K}\!N\!N$, considered to have quantum numbers $I=1/2$ and $J^{P}=0^{-}$, is essential for understanding the system’s fundamental features. We conducted the J-PARC E15 experiment to search for the $\bar{K}\!N\!N$ and reported the observation of the $I_{3}=+1/2~{}\bar{K}\!N\!N$ state, also known as ”$K^{-}pp$”, produced in the ${}^{3}{\rm He}(K^{-},~{}n)$ reactions and decaying into the non-mesonic $\Lambda p$ decay channel [1, 2]. The binding energy and decay width of $\bar{K}\!N\!N^{(I_{3}=+1/2)}$ were determined to be approximately $B_{\bar{K}\!N\!N}\sim 40~{}{\rm MeV}$ and $\Gamma_{\bar{K}\!N\!N}\sim 100~{}{\rm MeV}$, respectively, as the Breit-Wigner parameters.
In theoretical studies, the binding energy and decay width of $\bar{K}\!N\!N$ are calculated to be in the range of $10-50~{}{\rm MeV}$ and $40-70~{}{\rm MeV}$, respectively, using the chiral-SU(3) based potential [3, 4, 5, 6, 7, 8, 9]. Theoretical studies with phenomenological potentials reported significantly larger binding energies by $\sim 20~{}{\rm MeV}$ and similar decay widths [5, 6, 7, 8, 9]. Therefore, the obtained $B_{\bar{K}\!N\!N}$ is within the range of the theoretical calculations. However, $\Gamma_{\bar{K}\!N\!N}$ is larger than those predictions. Since the $\Lambda(1405)$ resonance is considered as the $\bar{K}\!N$ quasi-bound state, the dominant decay branch of $\bar{K}\!N\!N$ is naively expected to be $\bar{K}\!N\!N\to\pi\Sigma N$, similar to the $\Lambda(1405)\to\pi\Sigma$ decay (100%). Then $\Gamma_{\bar{K}\!N\!N}$ would be expected to be similar to the decay width of $\Lambda(1405)$. However, the obtained $\Gamma_{\bar{K}\!N\!N}$ is approximately twice as large as $\Gamma_{\Lambda(1405)}$.
As a reason for the large $\Gamma_{\bar{K}\!N\!N}$, two reasonable scenarios exist. One is that the $\bar{K}\!N\!N$ may decay into non-mesonic $Y\!N$ channels with similar strength to the $\pi\Sigma N$ channel. This non-mesonic decay occurs via $\bar{K}$ absorption by two nucleons in the nucleus, which most theoretical calculations did not consider. Another possibility is that the $I_{\bar{K}\!N}=1$ interaction may play an important role, similar to the $I_{\bar{K}\!N}=0$ interaction, at least in the absorption channels, i.e. $\bar{K}\!N\!N\to[\pi Y]_{I=1}N$, in contrast to the theoretical predictions [10, 11]. These two decay modes cannot occur in the decay of $\Lambda(1405)$ and newly open up in the $\bar{K}\!N\!N$ three-body system. Therefore, studying the mesonic decay modes is essential to clarify why $\Gamma_{\bar{K}\!N\!N}$ is much larger than $\Lambda(1405)$.
To investigate the mesonic decay branches of $\bar{K}\!N\!N$, we analyzed $K^{-}+^{3}{\rm He}$ reactions resulting in $\pi Y\!N\!N$ mesonic final states using data from the J-PARC E15 experiment. The measured final states in this study are $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, and $\pi^{\mp}\Sigma^{\pm}pn$. As described in Ref. [2], the $\bar{K}\!N\!N$ state is produced through the $\bar{K}$ and $N$ substitution via a nucleon knockout reaction on ${{}^{3}\rm He}$. Therefore, we focus on the $\pi Y\!N\!N$ channels that originate from the reaction processes:
$K^{-}+N\to{\bar{K}}+N^{\prime},$ (1)
where $N^{\prime}$ is emitted in the forward direction. Subsequently, the backscattered intermediate $\bar{K}$ is absorbed by the two spectator nucleons as:
${\bar{K}}+NN\to\pi Y\!N,$ (2)
where $\bar{K}\!N\!N$ production could contribute. The relevant $K^{-}+{{}^{3}\rm He}$ reaction channels of interest can be summarized as:
$\{\pi^{-}\Lambda p\}+p^{\prime}\to[\pi^{-}(\pi^{-}p)\,p]_{\rm det}+p_{\rm miss},$ (3)
$\{\pi^{+}\Lambda n\}+n^{\prime}\to[\pi^{+}(\pi^{-}p)\,n]_{\rm det}+n_{\rm miss},$ (4)
$\{\pi^{\mp}\Sigma^{\pm}p\}+n^{\prime}\to[\pi^{\mp}(\pi^{\pm}n)\,p]_{\rm det}+n_{\rm miss},$ (5)
where $\{\pi Y\!N\}$ is the particle set to be investigated, and $N^{\prime}$ is the emitted nucleon in the primary reaction (Eq.( 1)). Each final state is reconstructed as described on the right side of the equations, where $[\pi(\pi N)N]_{\rm det}$ represents detected particles, and $N_{\rm miss}$ is an undetected particle identified using the missing mass method. The $\bar{K}\!N\!N$ signal is expected to be observed in the $[\pi Y\!N]$ invariant-mass spectra for the $I_{3}=-1/2$ state in Eq.(3), and the $I_{3}=+1/2$ state in Eqs.(4) and (5). Following the analysis in Ref. [2], we performed a two-dimensional analysis on the $[\pi Y\!N]$ invariant-mass and the momentum transfer to the $[\pi Y\!N]$ system to isolate the signal of $\bar{K}\!N\!N$.
These reaction processes may couple to $Y^{*}$ resonances. In particular, two $Y^{*}$ resonances below the $\bar{K}\!N$ mass threshold, $\Lambda(1405)$ and $\Sigma(1385)$, would have an important contribution to the $\bar{K}$ absorption inside $\bar{K}\!N\!N$. Therefore, we analyzed the invariant-mass spectra of the $[\pi Y]$ subsystem to study the contributions of these $Y^{*}$ resonances.
This article is organized as follows. In Sec. II, we briefly introduce the J-PARC E15 experiment and describe the essential components of the detector system. In Sec. III, we present the analysis procedure used to select the $\pi Y\!N\!N$ channels. The obtained spectra will be shown in those results in Sec. IV. In Sec. V, we perform a model fitting using the functions introduced in Ref. [2]. Specifically, we discuss the $\bar{K}\!N\!N$ production and the quasi-free $\bar{K}$ absorption processes. In the fitting, we considered a background originating from the misidentification of $N^{\prime}$ and $N$ in Eqs.(1) and (2), respectively. Finally, we compare the cross sections of $K^{-}+{{}^{3}\rm He}$ reactions leading to $\pi Y\!N+N^{\prime}$ channels with the non-mesonic $\Lambda p+n$ channel.
II Experiment
We briefly overview the experimental setup of J-PARC E15, focusing on the key apparatus used to analyze the $\pi Y\!N\!N$ channels. The experimental configuration has been described in detail in the references [12, 13, 14]. We took the current data set in parallel to the one used for the $K^{-}+\,^{3}{\rm He}\to\Lambda p+n$ analysis [2] with the same trigger condition. In the experiment, a focused $K^{-}$ beam interacted with a liquid 3He target. The particles produced in the $K^{-}+{{}^{3}\rm He}$ reaction were detected by a cylindrical detector system (CDS) surrounding the target. The CDS consists of a cylindrical drift chamber (CDC) and a cylindrical hodoscope counter made of a plastic scintillator (CDH).
The CDS operates in a magnetic field of $0.7~{}{\rm T}$ along the beam axis ($z$-axis). The tracking information provided by the CDC allows for the analysis of charged particle momenta. We determined the starting point of the charged particle or the reaction point from the CDC track and $K^{-}$ beam track by searching for the closest point of the two tracks. Additionally, particle identification for each track is performed using momentum and time-of-flight information obtained from the CDH.
We employed the CDH for the neutron detection, which consists of plastic scintillators measuring 30 mm in thickness and 790 mm in length. The inner surface of the CDH is positioned at a radius of 544 mm from the beam axis. Using a Monte Carlo simulation based on Geant4, we estimated the typical neutron detection efficiency to be 3–9% (at neutron momenta of 1–0.2 ${\rm GeV}/c$). The neutron momentum was determined using the time-of-flight information and neutron flight length obtained from the reaction point, and the $z$-position along the CDH deduced by the time difference between two PMTs at both ends.
III Analysis
We focused on the following four channels: ${\pi^{-}\Lambda p}+p^{\prime}$ (Eq.(3)), ${\pi^{+}\Lambda n}+n^{\prime}$ (Eq.(4)), and ${\pi^{\mp}\Sigma^{\pm}p}+n^{\prime}$ (Eq.( 5)). Among these four channels, only the ${\pi^{-}\Lambda p}+p^{\prime}$ channel can be measured without detecting a neutron; namely, we required the detection of $(pp\pi^{-}\pi^{-})$ in the CDS. In the other three channels, neutron detection is necessary, along with detecting the remaining three charged particles ($p\pi^{+}\pi^{-}$).
The CDS detects the ${\pi Y\!N}$ particles in each final state. The undetected particle $N^{\prime}$ is identified through the missing mass method utilizing the measured $K^{-}$ beam and the information from the ${\pi Y\!N}$ particles. Other ${\pi Y\!N}+N^{\prime}$ channels, including multi-neutrons or $\gamma$ in the final decay products, were omitted from the present analysis since it is impossible to identify the final states without ambiguity.
III.1 Neutron analysis
We applied the following criteria to identify neutrons. Firstly, we required no CDC track on the fired CDH segment and both sides to eliminate fake neutron signals from charged particles hitting the material close to the CDH segment. Secondly, we required the energy deposited by the CDH-hit to be greater than $2~{}{\rm MeV}ee$ (electron-equivalent) to remove accidental background recorded in the CDH. Finally, $1/\beta$ was calculated for the CDH-hit using the time-of-flight information, and events with $1.51/\beta<$ 1.5 were rejected to eliminate $\gamma$-ray background originating from the $K^{-}+{{{}^{3}}{\rm He}}$ reaction which peaks at $1/\beta\approx 1$.
Figure 1: Mass distributions of the missing particle $X$ in the ($pn\pi^{+}\pi^{-}+X$) event selection. (a) shows the distribution before background subtraction, and (b) shows the distribution after subtracting the background.
Even after implementing the background rejection procedures described above, our data still contains numerous neutron-like fake signals in the neutron’s missing mass spectrum. Figure 1-(a) illustrates the distribution of missing mass ($m_{X}$) in the $K^{-}+{{{}^{3}}{\rm He}}\to pn\pi^{+}\pi^{-}+X$ events. The figure shows that the signal-to-noise ratio is approximately one-to-one, necessitating a thorough evaluation of the background effect for detailed analysis.
The fake neutron signals are generated through multiple reactions or scattering, reach the CDH, and are identified as neutral particles. Therefore, we evaluated the impact of these fake signals by simulating random CDH hits using an event-mixing method. In this method, we disconnected CDH-hit information from other charged particle hit information and randomly recombined the two data sets. The gray hatched curve in the figure represents the result of applying this event-mixing method. We determined a scaling factor for the background distribution by fitting the $m_{X}$-distribution below $0.8~{}{\rm GeV}/c^{2}$, where the data should only include neutron contamination events.
Figure 1-(b) shows that the background subtraction is successful in the $m_{X}=m_{n}$ region, and the $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$ final state remains clearly as a peak in the missing mass spectrum. This background subtraction procedure also agrees in the higher mass region up to 1.2 GeV/$c^{2}$, including the $m_{X}\sim m_{\Lambda}$ region. In fact, $\Lambda$ missing events, specifically $[pn\pi^{+}\pi^{-}]+\Lambda_{\rm miss}$ events, also remain clearly visible. This agreement demonstrates the validity of the fake-neutron subtraction procedure. We evaluated the fitting error of the scaling factor, as indicated by the gray box in the figure, and considered this error to be a systematic error in the subsequent analysis.
III.2 Selection of final state
For all the selected events, $[pp\pi^{-}\pi^{-}]+p_{\rm miss}$ and $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$, we applied a common procedure to identify the final states: $[\pi^{-}\Lambda p]+p_{\rm miss}$ (Eq.(3)), $[\pi^{+}\Lambda n]+n_{\rm miss}$ (Eq.(4)), and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ (Eq.(5)). We used a log-likelihood method, which was introduced in Ref. [2], to determine whether or not the event kinematics and topology were consistent with the reactions described in Eqs.(3), (4), and (5). To assess the kinematic information, we performed a kinematic fit to evaluate how well the event kinematics matched the final state of interest within the detector’s resolution. For geometrical information, we utilized the distances of the closest approach (DCA) between two charged particles. These distances should be zero within the resolution of the detector.
We defined the log-likelihood $L$ using probability density functions, $p(x)$, for the $\chi^{2}$ of the kinematical fit and DCAs, as follows:
$L=-\ln\left(p(\chi^{2}_{\text{kfit}})\times\prod_{i}^{N_{\text{DCA}}}p(\text{DCA}_{i})\right),$ (6)
where $N_{\text{DCA}}$ is the number of DCAs considered. $p(\chi^{2}_{\text{kfit}})$ represents the probability function of the $\chi^{2}$ with ${\text{NDF}}_{\text{kfit}}=2$. $p(\text{DCA})$ is a Gaussian function with mean $\mu=0$ and standard deviation $\sigma$, evaluated through Monte Carlo simulation with realistic detector resolutions. All $p(x)$ functions are normalized such that their maximum values are equal to one, making $L=0$ at the most probable point.
Multiple-particle combinations could be generated by hyperon decays or $\bar{K}^{0}\to\pi^{+}\pi^{-}$ decay. We examined all $L$ with possible particle combinations for each event to isolate which particle pairs are the decay products. Then, we determined that the reaction process of the events resulted in the minimum $L$. Figure 2 illustrates the distributions of the $L/{\rm NDF}$ for each identified process. As shown in the figure, the $L/{\rm NDF}$ distributions are quite similar, with the most frequent $L/{\rm NDF}$ values clustering around $L/{\rm NDF}\sim 0.5$. The yields decrease as $L/{\rm NDF}$ becomes larger. To select the final event sets, we accepted events with $L/{\rm NDF}<2$, and the threshold is indicated by the vertical red line in the figure.
Figure 2: Log-likelihood distributions for the event selections of (a) $\pi^{-}\Lambda pp$, (b) $\pi^{+}\Lambda nn$, (c) $\pi^{-}\Sigma^{+}pn$, and (d) $\pi^{+}\Sigma^{-}pn$. The red vertical line indicates the threshold $L/{\rm NDF}<2$, and events falling below this line are selected as the final event samples.
Due to the finite resolution, the correct particle combination for the reaction process is not always associated with the minimum $L$. There are three significant sources of contamination: I) combinatorial effects (improper particle pairing in hyperon decay, including $K^{0}$ contamination), II) $\Sigma^{0}$ contamination (resulting from a missing $\gamma$ in the $\Sigma^{0}$ decay), and III) contamination arising from the $\pi\pi Y\!N\!N$ final states. We assessed the contamination yields in each channel through simulation, as shown in the hatched region in the figure. The relative background yields to the total events in the $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ event sets were found to be $({\rm I}+{\rm II}+{\rm III})\sim(0+15+0)\%$, $(4+12+6)\%$, and $(5+0+6)\%$, respectively.
To demonstrate the validity of the event selection, we present the event distributions of the invariant mass of the selected $\pi N$ pairs from the hyperon decay ($m_{\pi N}$) and the missing mass of $\pi Y\!N$ ($m_{X}$) in Fig. 3. The top panel of the figure pair displays events with $L<100$, while the bottom panel represents events after the selection. The signals corresponding to each $[\pi Y\!N]+N_{\rm miss}$ final state are clearly observed as concentrated events around $(m_{X},~{}m_{\pi N})\sim(m_{N},~{}m_{Y})$, where $m_{N}$ and $m_{Y}$ represent the intrinsic masses of the nucleon and $\Lambda$ or $\Sigma$, respectively.
Figure 3: Two-dimensional distributions of the masses of the hyperon candidate ($m_{N\pi}$) and the missing particle ($m_{X}$). Panels (a), (b), (c), and (d) correspond to the event selections of $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, $\pi^{-}\Sigma^{+}pn$, and $\pi^{+}\Sigma^{-}pn$, respectively. The top figures display all events with $L<100$, while the bottom figures show the selected events with $L<2~{}{\rm NDF}$.
III.3 Acceptance correction
In the measurement, the acceptance of an event is determined by the four particles to be measured, $[\pi Y\!N]\to[\pi(\pi_{Y}N_{Y})N]$. We consider the event kinematics as a cascading two-body decay scheme illustrated in Fig. 4.
Figure 4: Considered kinematics for the $K^{-}+^{3}{\rm He}\to[\pi Y\!N]+N_{\rm miss}$ reaction. The figure illustrates the six kinematical parameters (indicated by red) that we considered for the acceptance correction.
In the reaction process of interest, the strange quark is carried by an intermediate $\bar{K}$, as discussed in Sec. I. Thus, the following equations can describe the process:
$\begin{split}K^{-}+~{}^{3}{\rm He}\to&[\bar{K}\!N\!N]+N_{\rm miss},\\ &[\bar{K}\!N\!N]\to[\bar{K}N]+[N].\end{split}$ (7)
This reaction corresponds to the two-body decays:
$\begin{split}K^{-}+~{}^{3}{\rm He}\to&[\pi Y\!N]+N_{\rm miss},\\ &[\pi Y\!N]\to[\pi Y]+[N],\end{split}$ (8)
as shown in the figure. The kinematics of each vertex point are defined by three parameters: the polar angle ($\theta$) and azimuthal angle ($\phi$) of the daughters in the rest frame of the parent system and the parent mass. The masses of the initial $K^{-}+~{}^{3}{\rm He}$ system and the final $\pi_{Y}+N_{Y}$ system are constrained to be $\sqrt{s}$ and $m_{Y}$, respectively. Therefore, there are ten independent kinematical parameters: the masses of $[\pi Y\!N]$ and $[\pi Y]$, $\theta$ and $\phi$ of $N_{\rm miss}$, $N$, $\pi$, and $\pi_{Y}$.
Since the azimuthal angles ($\phi$) do not affect the kinematics of the successive process, we can accurately evaluate the acceptance by integrating out $\phi$, assuming a uniform distribution. However, the polar angles ($\theta$) cannot be integrated. Consequently, we must consider the six kinematic parameters in the figure for the acceptance correction.
We conducted a Monte Carlo simulation using Geant4 to obtain the acceptance map. In the simulation, the $K^{-}+{{}^{3}\rm He}\to\pi Y\!N\!N$ reaction was generated according to the four-body phase space distribution. Figure 5 displays the distributions of the five kinematical parameters mentioned in Fig. 4 for both the experimental data and the simulation.
Figure 5: Distributions of the five kinematical parameters considered in the acceptance correction for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel. The black and blue distributions represent the data, and the four-body phase space distribution, including the acceptance. The simulated distributions were scaled to have the same integral as the data.
If the reaction is governed by the phase space only without any additional physics effects beyond the density of states of the final state, we would expect the data (black) and the simulation (blue) to coincide.
As shown in Fig. 5 (a) to (c), the data deviates from the simulation, but there is a good agreement in (d) to (f). Hence, we applied the acceptance correction event-by-event basis for these three parameters while integrating and correcting the remaining two using the average acceptance. Furthermore, we used $q_{[\pi Y\!N]}$ instead of $\theta_{N_{\rm miss}}$ since our interest lies in the momentum transfer to the $[\pi Y\!N]$ system rather than the specific formation angle $\theta_{N_{\rm miss}}$. Consequently, we constructed a three-dimensional acceptance map, $\varepsilon(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]},~{}m_{[\pi Y]})$, which encompasses the geometric acceptance of the CDS, detection efficiencies, and analysis efficiencies.
We are specifically interested in three distributions, namely $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$, for each $[\pi Y\!N]+N_{\rm miss}$ channel to investigate the production of $\bar{K}\!N\!N$ or $Y^{*}$ resonances. To obtain the differential cross sections of these three parameters, we initially calculated the triple differential cross section $d^{3}\sigma/dm\,dq\,dm_{[\pi Y]}$, where $m$ represents $m_{[\pi Y\!N]}$, $q$ represents $q_{[\pi Y\!N]}$. The expression is given by:
$\frac{d^{3}\sigma}{dm\,dq\,dm_{[\pi Y]}}=\frac{N(m,\,q,\,m_{[\pi Y]})-N_{\rm fake}(m,q,m_{[\pi Y]})}{\mathcal{L}\,\Delta m\,\Delta q\,\Delta m_{[\pi Y]}\,\varepsilon(m,q,m_{[\pi Y]})},$ (9)
where $\mathcal{L}=2.89\pm 0.01~{}{\rm nb}^{-1}$ represents the integrated luminosity, $N(m,\,q,\,m_{[\pi Y]})$ is the number of events in the ($m,\,q,\,m_{[\pi Y]}$) bin with bin widths $\Delta m$, $\Delta q$, and $\Delta m_{[\pi Y]}$. $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})$ denotes the expected fake-neutron background yield in the real data, as described in Sec. III.2 (note that $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})=0$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel, since there are no neutrons in this channel).
We omitted low-efficiency region $\varepsilon(m,q,m_{[\pi Y]})<0.1\%$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$, and $\varepsilon(m,q,m_{[\pi Y]})<0.015\%$ for other channels. Finally, we obtained differential cross-sections of interest by integrating the triple-differential cross-sections.
IV Results
Figure 6: Acceptance-corrected $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions and their projections on the $m_{[\pi Y\!N]}$ axes are shown in the figure: (a) corresponds to $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) corresponds to $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) corresponds to $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) corresponds to $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. The events-concentrations along a kinematical line (Eq.(10)) are clearly shown, especially in (a). For the mass region $m_{\pi^{-}\Lambda p}>2.6~{}{\rm GeV}/c^{2}$, the cross-section in (a) is scaled by a factor of $1/5$ to show events-concentration corresponding to Eq.(11). The black dotted lines represent the kinematical limit of the reaction. The blue vertical dotted lines indicate the mass threshold of $m_{\bar{K}}+2m_{N}$. The red dotted curves represent the quasi-free kinematical line described in Eq.(10).
IV.1 ($m_{\pi Y\!N},~{}q_{\pi Y\!N})$ distributions
We conducted measurements of the two-dimensional $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions to study the production of $\bar{K}\!N\!N$ decaying into mesonic $[\pi Y\!N]$ channels. Figure 6 shows the obtained $(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]})$ distributions for each final state, along with projections onto the $m_{[\pi Y\!N]}$ axis.
An events-concentration at $(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]})\sim(2.4~{}{\rm GeV}/c^{2},~{}0.3~{}{\rm GeV}/c)$ was observed consistently in all $\pi Y\!N\!N$ channels. The origin of these events-concentrations can be understood as follows. The concentrations are along the kinematical line of the quasi-free $\bar{K}$ absorption by residual two nucleons given by
$m(q)=\sqrt{4m_{N}^{2}+m_{\bar{K}}^{2}+4m_{N}\sqrt{m_{\bar{K}}^{2}+q^{2}}},$ (10)
where $m_{N}$ and $m_{\bar{K}}$ are the intrinsic masses of the $N$ and $\bar{K}$, respectively. Equation (10) describes the invariant mass of $\bar{K}$ with momentum $q$ interacting with two residual nucleons, thereby representing the kinematics of a quasi-free absorption process where two residual nucleons absorb an on-shell $\bar{K}$ at rest. Given that the events are concentrated along this kinematical line, we can interpret this concentration as a manifestation of the quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) process, similar to the non-mesonic $[\Lambda p]+n_{\rm miss}$ channel [2].
If the $\bar{K}\!N\!N$ quasi-bound state observed in the $[\Lambda p]+n_{\rm miss}$ channel [2] decay into the mesonic channels, one can expect to observe spectral peaks in the invariant-mass spectra of each $[\pi Y\!N]$ channel below the kaon binding threshold, $m_{\bar{K}\!N\!N}=m_{\bar{K}}+2m_{N}$ (indicated by the blue dotted line in the figure). However, clear peak structures were not found in these spectra. The absence of clear peak structures can be understood by considering the density of states in the final state, which is governed by the four-body phase space of $[\pi Y\!N]+N_{\rm miss}$. The phase space volume below the $m_{\bar{K}\!N\!N}$ threshold becomes very small since the mass region is close to the $\pi Y\!N$ threshold, and the four-body phase space opens from the $\pi Y\!N$ threshold only in the second order. As a result, the invariant-mass spectra of $[\pi Y\!N]$ in this region are significantly modified due to the limited phase space near the threshold. This spectral modification leads to the disappearance of the prominent peak structure. The density of states in the mesonic final states and the resulting spectral modification will be described in more detail in Sec. V.
[…]
$K^{-}+(pn)\to\pi^{-}\Lambda p.$ (11)
In this process, the $\bar{K}$ reacts only with a deuteron-like momentum correlated $pn$-cluster in ${{}^{3}\rm He}$. The other proton in ${}^{3}{\rm He}$ acts as a spectator with a Fermi momentum, serving as the missing particle ($p_{\rm miss}$). We plotted the momentum distribution of $p_{\rm miss}$ in Fig. 7 to investigate this interpretation. As shown in the figure, the momentum distribution of $p_{\rm miss}$ is consistent with our spectator picture, exhibiting a low momentum component in agreement with Fermi motion in ${{}^{3}\rm He}$ [15].
Figure 7: Momentum distribution of the missing proton ($p_{\rm miss}$) in the $(\pi^{-}\Lambda p+p_{\rm miss})$ channel. The gray dotted and solid lines represent the Fermi momentum distribution in ${{}^{3}\rm He}$ [15].
In contrast, the same events-concentration was not observed in the $[\pi^{+}\Lambda n]+n_{\rm miss}$ channels, as shown in Fig. 6-(b). This result suggests that the $K^{-}$ beam is less likely to be absorbed by a ($pp$) pair compared to a ($pn$) pair, whose ratio ($pp$/$pn$) is less than $1/10$. This result is consistent with the absorption of stopped-$\pi^{-}$ on ${{}^{3}\rm He}~{}$[16].
IV.2 $m_{\pi Y}$ distributions
To investigate the role of $Y^{*}$-resonances, if the events coupling to the $Y^{*}$-poles, we plotted the invariant-mass distribution of the $[\pi Y]$ sub-system, specifically $d\sigma/dm_{[\pi Y]}$, as shown in Fig. 8. For the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel, we have excluded events originating from the direct-2NA process by removing events with $p_{p_{\rm miss}}^{\rm lab}<0.3~{}{\rm GeV}/c$. This criterion facilitates a more straightforward comparison with other channels.
The figures illustrate that the overall shapes of the $m_{\pi Y}$ distributions are similar, exhibiting a clear two-body decay of $Y^{*}\to\pi Y$ over a broad non-resonant background having similar yields. Regarding the $[\pi^{\pm}\Lambda]$ distributions, the positions of the peaks are consistent with those of $\Sigma(1385)^{\pm}$. This consistency can be attributed to the fixed isospin of $\pi^{\pm}\Lambda$, resulting in a natural coupling to $\Sigma(1385)^{\pm}$. In contrast, the observed peaks in the $[\pi^{\mp}\Sigma^{\pm}]$ distributions are slightly higher than $1.4~{}{\rm GeV}/c^{2}$. Hence, it can be inferred that these peaks primarily arise from the coupling to the $\Lambda(1405)$ pole. In this channel, the isospin can be either zero or one. However, the contribution from $\Sigma(1385)^{0}\to\pi^{\mp}\Sigma^{\pm}$ is expected to be negligible. This is because the decay branching ratio of $\Sigma(1385)$, $Br(\pi\Sigma)/Br(\pi\Lambda)$, is approximately 13%, and if the isospin symmetry holds between $\Sigma(1385)^{0}$ and $\Sigma(1385)^{\pm}$, their production yields should be of the same order. No higher mass $Y^{*}$ resonances were observed in the $[\pi Y]$ distributions.
In addition to the study of the $[\pi Y]$ subsystem, we investigated the $[\pi N]$ subsystem to identify if there are any spectral anomalies associated with $\Delta$ or $N^{*}$ resonances. The reactions under consideration were:
$K^{-}+{{}^{3}\rm He}\to Y\Delta N~{}{\rm or}~{}YN^{*}N\to[Y(\pi N)]N_{\rm miss}.$ (12)
The resulting $d\sigma/dm_{\pi N}$ spectra are presented in Fig. 9. In contrast to the $m_{\pi Y}$ distributions, the $m_{\pi N}$ distributions did not exhibit any clear peak or structure.
The obtained results indicate that the dominant processes involve a backward low-momentum $\bar{K}$ in the intermediate state of the reaction via the nucleon knockout reaction $K^{-}N\to\bar{K}N^{\prime}$. This $\bar{K}$ could efficiently couple with $\Sigma(1385)$ or $\Lambda(1405)$ via $\bar{K}+NN\to Y^{*}N\to\pi Y\!N$.
Figure 8: Acceptance-corrected $m_{\pi Y}$ distributions for each $(\pi YNN)$ final state. The blue dotted line represents the mass threshold of $m_{\bar{K}}+m_{N}$ ($m_{\bar{K}N}$).
Figure 9: Acceptance-corrected $m_{\pi N}$ distributions for each $(\pi Y\!N\!N)$ final state. The black dotted line represents the mass threshold of $m_{\pi}+m_{N}$.
V Discussion
As previously described, the overall structures of the mesonic $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions (see Fig. 6) are similar to the non-mesonic $(m_{[\Lambda p]},q_{[\Lambda p]})$ distribution [2], assuming the phase-space suppression of the $\bar{K}\!N\!N$ peak in mesonic decay channels. The $m_{[\pi Y]}$ spectra presented in Fig. 8 consist of a prominent peak associated with the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, as well as a broad non-resonant contribution.
Therefore, we examined if the event distribution of $\pi Y\!N\!N$ final state can be described essentially in the same model functions applied for $\Lambda pn$ final state, which is composed of $\bar{K}\!N\!N$ formation and ${\rm QF}_{\bar{K}-{\rm abs}}$ as it is described in Ref. [2]. Thus, we assumed that the same event distribution for the intermediate $(\bar{K}+NN)$ system is formed over $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]})$ kinematical plane as it is the case for $\Lambda pn$ final state, and the intermediate $(\bar{K}+NN)$ decays to $\pi Y\!N$. To allow the $\pi Y$ subsystem to couple to the $Y^{*}$-poles in the decay process, we modified the model fitting function $F$ and the phase space accordingly. Therefore, the present model fitting function becomes as follows:
$\begin{split}F(m,q,m_{\pi Y})=\rho(m,q,m_{\pi Y})\,\varepsilon(m,q,m_{\pi Y})\\ \times\sum_{j}A_{j}f_{j}(m,\,q)\big{(}1+A_{Y^{*}}\,f_{BW}(m_{\pi Y})\big{)}\\ +{\rm BG}(m,\,q,\,m_{\pi Y}),\end{split}$ (13)
where $\rho$ represents the four-body phase-space for $\pi Y\!N\!N$, $\varepsilon$ denotes the acceptance, $f_{BW}$ is a Breit-Wigner function used to describe the $\Sigma(1385)$ or $\Lambda(1405)$ resonance, $A_{Y^{*}}$ is the coupling strength to the $Y^{*}$ pole, and $f_{j}$ and $A_{j}$ represent the spectral function and strength of each physical process ($j=\bar{K}\!N\!N$ and ${\rm QF}_{\bar{K}-{\rm abs}}$), respectively. The parameters $m_{Y^{*}}$ and $\Gamma_{Y^{*}}$ of $f_{BW}$ were fixed to the values provided by the Particle Data Group (PDG) for the case of $\Sigma(1385)$ [17], while we allowed to vary these values for the case of $\Lambda(1405)$. In the fitting procedure, we ignored any interference between reaction processes. A resolution function, which smears $f_{i}$ and $f_{BW}$, was evaluated by the Monte Carlo simulation and convoluted to Eq.(13).
The $\bar{K}\!N\!N$ spectral function $f_{K}$ is defined as follows:
$\begin{split}f_{K}(m_{\pi Y\!N},q_{\pi Y\!N})=\frac{(\Gamma_{K}/2)^{2}}{(m_{\pi Y\!N}-M_{K})^{2}+(\Gamma_{K}/2)^{2}}\\ \times A_{0}^{K}\exp\left(-\frac{q_{\pi Y\!N}^{2}}{Q_{K}^{2}}\right),\end{split}$ (14)
where $M_{K}$, $\Gamma_{K}$, $Q_{K}$, and $A_{0}^{K}$ are parameters. The ${\rm QF}_{\bar{K}-{\rm abs}}$ spectral function $f_{F}$ is a Gaussian distribution along the $M_{F}(q_{\pi Y\!N})$ (Eq.(10)) given by:
$\begin{split}f_{F}(m_{\pi Y\!N},q_{\pi Y\!N})=\exp\left[-\frac{\left(m_{\pi Y\!N}-M_{F}(q_{\pi Y\!N})\right)^{2}}{\left(\sigma_{0}+\sigma_{2},q_{\pi Y\!N}^{2}\right)^{2}}\right]\\ \times\left[A_{0}^{F}\exp\left(-\frac{q_{\pi Y\!N}^{2}}{Q_{F}^{2}}\right)+A_{1}^{F}+A_{2}^{F}\exp\left(\frac{m_{\pi Y\!N}}{m_{0}}+\frac{q_{\pi Y\!N}}{q_{0}}\right)\right],\end{split}$ (15)
where $\sigma_{0}$, $\sigma_{2}$, $Q_{F}$, $m_{0}$, $q_{0}$, $A_{0}^{F}$, $A_{1}^{F}$, and $A_{2}^{F}$ are parameters. In the current fitting, we fixed the parameters of $f_{K}$ and $f_{F}$ to those obtained in Ref. [2].
To illustrate how the phase space difference can modify spectral functions, we present Fig. 10, which shows two-dimensional model fitting functions for the $\Lambda p$ and $[\pi Y\!N]$ spectra. The model functions for the $\bar{K}\!N\!N$ production process in each final state are shown in Fig. 10-(a1-3). In the $[\Lambda p]+n$ final state (Fig. 10-(a1)), the reaction threshold ($m_{\Lambda}+m_{N}$) is significantly smaller than $M_{K}$ in Eq.(14), resulting in a nearly flat density-of-state for $[\Lambda p]$ around the $\bar{K}\!N\!N$ peak. On the other hand, in the $[\pi\Lambda N]+N$ (Fig. 10-a2) and $[\pi^{\mp}\Sigma^{\pm}p]+n$ final states (Fig. 10-a3), the reaction threshold ($m_{\pi}+m_{Y}+m_{N}$) approach to $M_{K}$, resulting in substantial suppression of the spectrum due to the diminishing density-of-state at the lighter mass side of the peak structure, so that the maximum of the spectrum appears to shift towards the heavier side compared to $M_{K}$ value. Regarding the ${\rm QF}_{\bar{K}-{\rm abs}}$ process, the spectral modifications in the $[\pi Y\!N]+N$ channels are relatively minor compared to those observed in the $\bar{K}\!N\!N$ case, as illustrated in Fig. 10-(b1-3). This minor modification in the ${\rm QF}_{\bar{K}-{\rm abs}}$ process is primarily due to the energy region being further separated from the reaction threshold.
Figure 10: Fit functions for (a1, a2, a3) the $\bar{K}\!N\!N$ production process and (b1, b2, b3) the quasi-free kaon absorption on the two-nucleon $QF_{\bar{K}-abs}$ process. (a1) and (b1) display spectral functions using the three-body phase space of $\Lambda pn$. (a2) and (b2) correspond to spectral functions using the four-body phase space of $\pi\Lambda NN$. (a3) and (b3) represent spectral functions using the four-body phase space of $\pi\Sigma pn$. The $z$ scale is logarithmic. The black, blue, and red dotted lines correspond to the same lines as in Fig. 6.
This analysis assumes that the detected particles, $\pi Y\!N$, originate from the $(\bar{K}+NN)$ intermediate state. However, it is possible to detect knocked-out $N^{\prime}$ in the primary $K^{-}N\to\bar{K}N^{\prime}$ reaction without detecting $N$ from the intermediate state. In such cases, the $(m,q)$ values of the $[\pi Y]N^{\prime}$ system (rather than $[\pi Y\!N]$) are distributed widely across the kinematically allowed region. To incorporate this scenario, we generated expected spectral functions through simulations for each relevant process and summed them up to account for the background, ${\rm BG}(m,~{}q,~{}m_{\pi Y})$. There is background contamination resulting from the misidentification of the final state, as described in Sec. III.2. In the present fitting, we omitted this background but considered its effect on evaluating the cross-section a systematic error.
Due to the limited statistics over the limited acceptance on three-dimensional kinematical space of $(m_{[\pi Y\!N]},\,q_{[\pi Y\!N]},\,m_{[\pi Y]})$, a multi-dimensional fitting nor the application of the coupled channel based sophisticated fitting function (to account for the threshold effects) are both not feasible in this analysis. Instead, we performed a simultaneous fitting on three one-dimensional spectra along the $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$ axes. The two charged modes, $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$, were fitted together by using common values of $M_{\Lambda(1405)}$ and $\Gamma_{\Lambda(1405)}$. The fitting results are shown in Fig. 11, and the obtained parameters are summarized in Table 1. The chi-squares and the number of degrees of freedom for $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ were $\chi^{2}/NDF=275.5/118$, $76.3/85$, and $114.9/136$, respectively. The figures demonstrate that the proposed model functions provide a good overall data description.
Figure 11: The fitted distributions and the results of the fitting for the following channels: (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. “BS”, “QF”, and “BG” represent the $\bar{K}\!N\!N$ production, the ${\rm QF}_{\bar{K}-{\rm abs}}$, and the background processes, respectively.
Table 1: Parameters obtained by the fit.
$A_{i}$ for $(\pi^{-}\Lambda p)$ channel
$A_{K}$ $1140.5\pm 112.2$
$A_{F}$ $1800.6\pm 77.0$
$A_{Y^{*}}$ $1.4\pm 0.1$
$A_{i}$ for $(\pi^{+}\Lambda n)$ channel
$A_{K}$ $2707.7\pm 467.3$
$A_{F}$ $1557.5\pm 114.5$
$A_{Y^{*}}$ $1.2\pm 0.4$
$A_{i}$ for $(\pi^{-}\Sigma^{+}p)$ channel
$A_{K}$ $2010.6\pm 57.3$
$A_{F}$ $1167.9\pm 138.8$
$A_{Y^{*}}$ $4.4\pm 0.6$
$A_{i}$ for $(\pi^{+}\Sigma^{-}p)$ channel
$A_{K}$ $2159.7\pm 150.8$
$A_{F}$ $2583.9\pm 108.1$
$A_{Y^{*}}$ $1.8\pm 0.2$
parameters for $\Lambda(1405)$
$M_{\Lambda(1405)}$ $1432\pm 5~{}{\rm MeV}/c^{2}$
$\Gamma_{\Lambda(1405)}$ $49\pm 3~{}{\rm MeV}/c^{2}$
In Fig. 11-(c) and (d), it can be observed that the $\Lambda(1405)$ contribution in the $m_{\pi^{\mp}\Sigma^{\pm}}$ distributions is well reproduced by a simple Breit-Wigner function. The obtained mass and width of $\Lambda(1405)$ are $1432\pm 5~{}{\rm MeV}/c^{2}$ and $49\pm 3~{}{\rm MeV}/c^{2}$, respectively. These values are close to the pole position deduced from the $d(K^{-},n)\pi\Sigma$ reactions, $(m,\Gamma)=(1418~{}{\rm MeV}/c^{2},52~{}{\rm MeV}/c^{2})$[18], as well as the theoretically predicted pole positions based on chiral unitary approaches, $(1421-1434~{}{\rm MeV}/c^{2},20-52~{}{\rm MeV}/c^{2})$ [19, 20, 21, 22, 23, 24].
We examined the validity of the global fit using $q_{[\pi Y\!N]}$-sliced mass spectra, as shown in Figs. 12 and 13. The model functions successfully reproduced the global structure of each sliced distribution except for the $m_{[\pi Y]}$ spectra near the $\bar{K}\!N$ threshold in the low momentum transfer region of $q<0.3~{}{\rm GeV}/c$ as shown in Fig.13. This discrepancy may come from the neglected channel coupling effect with the $\bar{K}\!N$ channel, i.e., the $\bar{K}N$ threshold cusp in the $I_{\bar{K}N}=1$ channel recently reported by the Belle collaboration [25].
Figure 12: Differential cross sections $d\sigma/dm_{\pi Y\!N}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fit result obtained from the analysis. In panel (a), the gray hatched box highlights a region outside the fitting range to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N\!N}$ threshold.
Figure 13: Differential cross sections $d\sigma/dm_{\pi Y}$ are shown in each $q$ interval for (a) $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$ channels. The colored histograms represent the fit result obtained from the analysis. In the case of (a), requirements of $m_{\pi^{-}\Lambda p}<2.6~{}{\rm GeV}/c^{2}$ and $p_{\rm miss}>0.3~{}{\rm GeV}/c$ are imposed to exclude the direct 2NA process. The blue dotted lines indicate the $m_{\bar{K}\!N}$ threshold.
The total cross sections of $\bar{K}\!N\!N$ decaying into each mesonic decay channel, $\sigma_{\bar{K}\!N\!N}^{tot}\,Br(\pi Y\!N)$, were evaluated by integrating the $\bar{K}\!N\!N$ fitting function over the entire kinematical region including above the $\bar{K}$ binding threshold, $m_{\bar{K}}+2m_{N}$ as summarised in Tab. 2. We also tabulated the branch below the $m_{\bar{K}}+2m_{N}$ threshold to exclude the energy region where the spectral shape of the fitting functions would become less reliable due to threshold effects. The values of non-mesonic decay modes, $\bar{K}\!N\!N_{I_{3}=+1/2}\to\Lambda p$ and $\Sigma^{0}p$, were taken from Ref. [2]. The values of the $\bar{K}\!N\!N\to\Sigma^{+}n$ and $\bar{K}\!N\!N\to\pi^{0}\Lambda p$ channels were evaluated as $Br(\Sigma^{+}n)=2Br(\Sigma^{0}p)$ and $Br(\pi^{0}\Lambda p)=1/2Br(\pi^{+}\Lambda n)$, respectively, based on the isospin symmetry. The total non-mesonic decay branch was then evaluated as the sum of the three non-mesonic channels, the errors of which are the linear sum of those three channels since the errors are correlated.
We obtained that the ratio of non-mesonic to mesonic decay branches is $\mathcal{O}(10)$ if we integrate all mass regions. The mesonic decay branch is still larger than the non-mesonic one, even if we integrate below the $\bar{K}$ binding threshold. The result suggests that the mesonic decay mode is the dominant decay branch of the $\bar{K}\!N\!N$, as expected from theoretical considerations.
The decay branches $Br(\pi^{+}\Lambda n)$ and $Br(\pi^{\mp}\Sigma^{\pm}p)$ were similar in magnitude in both integration ranges. The result suggests that the mesonic decay modes coupled to the $I_{\bar{K}\!N}=1$ channel play a significant role in the $\bar{K}\!N\!N$ decay process. Furthermore, this could explain why the decay width of the $\bar{K}\!N\!N$ is much broader than that of the $\Lambda(1405)$ resonance, as the mesonic decay modes with $I_{\bar{K}\!N}=1$ contribute to the broadening of the $\bar{K}\!N\!N$ decay width.
Table 2: Cross sections of each decay channel of $\bar{K}\!N\!N$. The first and second errors are statistical and systematic ones, respectively. The $\bar{K}$ binding thresholds for $I_{3}=\pm 1/2$ were considered as $m_{K^{-}}+2m_{p}$ and $m_{\bar{K}^{0}}+2m_{n}$, respectively. We estimated decay branches of $\Sigma^{+}n$ and $\pi^{0}\Lambda p$ channels by assuming the isospin symmetry.
$\sigma_{\bar{K}\!N\!N}^{tot}\,Br~{}(\mu{\rm b})$
Decay channel all-region below the $\bar{K}$ binding threshold
$\bar{K}\!N\!N_{I_{3}=+1/2}$
$\Lambda p$ $9.3\pm 0.8^{+1.4}_{-1.0}$ [2] $5.5\pm 0.5^{+0.8}_{-0.6}$
$\Sigma^{0}p$ $5.3\pm 0.4^{+0.8}_{-0.6}$ [2] $3.1\pm 0.2^{+0.5}_{-0.4}$
$\Sigma^{+}n^{(*)}$ ($=\Sigma^{0}p\times 2$) $10.6\pm 0.8^{+1.6}_{-1.2}$ $6.2\pm 0.4^{+1.0}_{-0.8}$
total non-mesonic $25.2\pm 2.0^{+3.8}_{-2.8}$ $14.8\pm 1.1^{+2.3}_{-1.8}$
$\pi^{0}\Lambda p^{(*)}$ ($=\pi^{+}\Lambda n\times 1/2$) $31\pm 5.5\pm 4.5$ $7.8\pm 1.4\pm 1.1$
$\pi^{0}\Sigma^{0}p$ NA NA
$\pi^{-}\Sigma^{+}p$ $110\pm 8\pm 8$ $9.4\pm 0.4\pm 0.7$
$\pi^{+}\Sigma^{-}p$ $38\pm 3\pm 3$ $3.2\pm 0.2\pm 0.2$
$\pi^{+}\Lambda n$ $62\pm 11\pm 9$ $15.5\pm 2.7\pm 2.1$
$\pi^{+}\Sigma^{0}n$ NA NA
$\pi^{0}\Sigma^{+}n$ NA NA
total mesonic $>241\pm 20\pm 17$ $>37.9\pm 4.1\pm 3.3$
$\bar{K}\!N\!N_{I_{3}=-1/2}$
$\pi^{-}\Lambda p$ $29\pm 3\pm 3$ $7.2\pm 0.6\pm 0.7$
VI Summary
We conducted measurements of the $K^{-}+{{}^{3}\rm He}$ reactions resulting in mesonic final states, namely $\pi^{\mp}\Sigma^{\pm}p+n^{\prime}$, $\pi^{+}\Lambda n+n^{\prime}$, and $\pi^{-}\Lambda p+p^{\prime}$, using an incident $K^{-}$ momentum of $1~{}{\rm GeV}/c$. To investigate the mesonic decay of the $\bar{K}\!N\!N$ state, we measured the two-dimensional distributions of the invariant mass and momentum transfer of the $\pi Y\!N$ systems. The observed two-dimensional distributions of the $\pi Y\!N$ systems were well described by the model functions, which included contributions from $\bar{K}\!N\!N$ production and quasi-free $\bar{K}$ absorption processes as a result of reactions given in Eqs.(1) and (2). These model functions are essentially the same as those introduced in Ref. [2] to explain the $K^{-}+{{}^{3}\rm He}\to\Lambda p+n^{\prime}$ reaction by incorporating appropriate phase-space distributions.
The fit suggests that the mesonic decay is the dominant decay branch of the $\bar{K}\!N\!N$. It was also suggested that the branching ratio of $\bar{K}\!N\!N\to\pi^{+}\Lambda n$ is similar to that of $\bar{K}\!N\!N\to\pi^{\mp}\Sigma^{\pm}p$, indicating that the mesonic $\bar{K}N$ absorption process in the $I_{\bar{K}N}=1$ channel plays an important role similar to the $I_{\bar{K}\!N}=0$ channel. The significant fraction of decay into the $I_{\bar{K}N}=1$ channel, specifically the $\pi\Lambda N$ decay, would be a major contributing factor to the broader decay width of the $\bar{K}\!N\!N$ resonance compared to that of the $\Lambda(1405)$. The presence of the $I_{\bar{K}\!N}=1$ cusp in the $\pi Y$ subsystem (Fig. 13) suggests that the real part of the $I_{\bar{K}\!N}=1$ interaction is also attractive, although it is not sufficiently strong to form a bound state in this channel.
The branching ratio of the $\bar{K}\!N\!N$ between the non-mesonic and mesonic channels provides valuable information regarding the size of the system. This is because the strength of mesonic decay is directly proportional to the nucleon density. In contrast, the strength of non-mesonic decay is proportional to the square of the nucleon density, as it requires simultaneous interaction of the antikaon with two nucleons. Based on the present study, the $Br(YN)/Br(\pi Y\!N)$ ratio is estimated to be approximately 1/10 obtained as integration over all mass regions. However, we should note that the ratio becomes smaller if the integration range is up to the $\bar{K}$ binding threshold. A more precise analysis, especially using a more realistic model function, is necessary to compare the experimental results with the theoretical calculations. This analysis would provide more information about the size of the system.
Further study of the multi-nucleon absorption of the $K^{-}$ beam could provide valuable insights into the clustering mechanism within the nucleus. The data show that the kaon’s direct two-nucleon absorption (2NA) process is predominantly observed in the ”$pn$” pair. Conversely, almost no reaction is observed in the ”$pp$” pair. This difference can intuitively indicate a stronger coupling or clustering of the ($pn$) pair compared to the ($pp$) pair in ${{}^{3}\rm He}$.
It is crucial to conduct new experiments with enhanced acceptance coverage and improved neutron detection efficiency to understand these issues better and investigate the production mechanism of kaonic nuclei. These advancements will enable us to gather more comprehensive data and provide valuable insights into the phenomena under investigation. By increasing the experimental capabilities, we can further explore the properties and behavior of kaonic nuclei, shed light on their production mechanisms, and contribute to a more thorough understanding of these fascinating systems [26].
Acknowledgements.
The authors are grateful to the staff members of J-PARC/KEK for their extensive efforts, especially in the stable operation of the facility. We are also grateful to the contributions of Professors D. Jido, T. Sekihara, Y. Akaishi, A. Dote, T. Harada, O. Morimatsu, J. Yamagata, and Dr. K. Suzuki. This work is partly supported by MEXT Grants-in-Aid 14102005, 17070007, 24105003, 26287057, 26800158, 17K05481, 18H05402, 20K04006, 21K13952, and 22H04917. Part of this work is supported by the EU STRONG-2020 project (Grant Agreement No. 824093) and by the EXOTICA project of the Minstero degli Affari Esteri e della Cooperazione Internazionale, PO22MO03.